2.42.10cm3 of a solution containing 2g of an acid (H2X) per dm3 neutralizes 28cm3 of solution containing 2.5g of potassium hydroxide per dm3. Calculate
i.The concentration of the acid in moles/dm3
ii.The molar mass of the acid H2X
iii.The value of X
iv.The percentage by mass of X
1
Expert's answer
2020-06-18T14:42:15-0400

The balanced equation for the reaction of H2X with potassium hydroxide is:

H2X + 2KOH "\rightarrow" K2X + 2H2O.

According to the stoichiometry coefficients, 1 mol of H2X reacts with 2 mol of KOH:

"n(H_2X) = \frac{n(KOH)}{2}" .

The molar concentration of KOH solution is:

"c = \frac{n}{V} = \frac{m}{MV} = \frac{2.5(g)}{56.11(g/mol)·1(dm^3)} = 0.0446" mol/dm3.

The number of the moles of potassium hydroxide reacted is the product of the concentration and the volume of its solution:

"n(KOH) = cV"

"n(KOH)= 0.0446(mol\/dm^3)\u00b728\u00b710^{-3} (dm^3)"

"n(KOH) = 1.25\u00b710^{-3}" mol.

Therefore, the number of the moles of the H2X acid is:

"n(H_2X) = \frac{1.25·10^{-3}}{2} = 0.62·10^{-3}" mol.

The molar concentration of the acid is the number of the moles divided by its volume:

"c(H_2X) = \frac{n}{V} = \frac{0.62·10^{-3}(mol)}{42.10·10^{-3}(dm^3)} = 0.0148" mol/dm3.

The molar mass of H2X can be calculated from its molar concentration and mass concentration (g/dm3):

"M = \frac{m}{V}·\frac{1}{c} = 2(g/dm^3)·\frac{1}{0.0148(mol/dm^3)} = 135" g/mol.

The molar mass of X is then:

"M_X = M(H_2X) - 2M_H = 135-2 = 133" g/mol.

Using the data given, X can't be assigned to any element in the periodic table.

The percentage by mass of X in the molecule H2X is:

"\frac{M(X)}{M(H_2X)}·100\% = \frac{133}{135}·100\% =98.5\%" .


Answer:

i.The concentration of the acid is 0.0148 mol/dm3

ii.The molar mass of the acid H2X is 135 g/mol

iii.The data given is not consistent and doesn't permit to find the value of X

iv.The percentage by mass of X is 98.5%